OpenStudy (anonymous):

what is the antiderivative of (x + 1)/(xˆ2 +2x +3) dx?

OpenStudy (anonymous):

firstly, notice that when we differentiate the denominator we get a multiple of the numerator

OpenStudy (anonymous):

secondly recall that the chain rule applied to ln(u) \[\frac{d}{dx}\ln(u) = \frac{1}{u} \times \frac{du}{dx}\]

OpenStudy (anonymous):

whenever you see the derivative of a f(x) multiplied by a function of f(x) this chain rule. in this case differentiating ln(x^2 + 2x +3) with respect to x gets us (2x+2)/(x^2 +2x +3)

OpenStudy (anonymous):

so we want half of that

OpenStudy (anonymous):

*think ,not "this"

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

simple take denominator as "t" or some othr variable differentiate n substitute. let t =(xˆ2 +2x +3) dt=(2x+2)dx dt/2=(x+1)dx so our integral becomes  integrate n resubsitute t in final answer..